# Heavy Traffic Diffusion Limit for a Closed Queueing Network with Single-Server and Infinite-Server Stations

### v2

@@ 1 Introduction
The model was motivated in part by ride-hailing systems such as Uber and Lyft. In that context, the single-server stations represent regions of a city where drivers await ride requests, while the infinite-server stations capture travel times between regions. After picking up a passenger, a driver enters a travel phase to reach the destination, and upon completion, becomes available again in a potentially new region. The two-level routing structure in our model naturally reflects such origin-destination dynamics, and the presence of multiple infinite-server stations allows for heterogeneity in travel times across different region pairs. While ride-hailing serves as a useful motivation, the model and results apply more broadly to systems in which tasks are queued for assignment or dispatch at specific locations, and then processed or delayed in parallel elsewhere. By establishing a rigorous diffusion approximation for such systems, this paper lays the theoretical groundwork for future research on their dynamic control and optimization.
@@ 1.1 Literature Review
This paper contributes to the literature on heavy traffic diffusion approximations for queueing networks, with a particular focus on closed systems involving infinite-server stations. Classical work by Iglehart [19], Borovkov [10], and Whitt and Glynn [27] establishes diffusion limits for open queueing networks with infinite-server queues. Other papers have studied the structure and properties of these limiting processes. For example, Glynn [15] establishes necessary and sufficient conditions under which the heavy traffic limit of the $GI/G/\infty$ queue is Markovian. Foundational work on heavy-traffic approximations for stochastic flow networks with bottlenecks includes Chen and Mandelbaum [12, 13], which develop both fluid and diffusion limits. Surveys such as Glynn [16] and Pang et al. [24] review various techniques for deriving diffusion approximations in queueing systems; see the references therein for a broader overview.
@@
In the context of closed queueing networks with infinite-server stations, several papers establish diffusion limits under heavy traffic conditions. Notably, Kogan et al. [22], Krichagina [23], and Kogan and Lipster [21] analyze models with many single-server stations and a single infinite-server station. These papers have a single probability routing vector and use martingale methods to prove convergence. Our paper supplements this line of research by allowing an arbitrary number of infinite-server stations and a two-level probabilistic routing structure. (This enables more realistic modeling of origin-destination dynamics and heterogeneous travel times, which are important to service systems such as ride-hailing.) Our analysis also differs methodologically: we use a continuous mapping approach and establish existence, uniqueness, and continuity of a nonlinear regulator mapping—results that are not readily implied by existing work. To the best of our knowledge, this is the first paper to rigorously establish a diffusion limit for a closed queueing network with both multiple infinite-server stations and a two-level routing structure. Other papers have studied the asymptotic behavior of the stationary distribution of customers (specifically, the partition function) in closed queueing networks with multiple infinite-server stations and a single infinite-server station, including Birman and Kogan [9] and Kogan [20].
@@
A closely related and important antecedent of our model is the one developed in Alwan et al. [1], which considers a ride-hailing system modeled as a closed queueing network with multiple single-server regions and a single infinite-server travel node. The objective is to maximize system profit by making dynamic pricing and dispatch control decisions. That paper formulates an approximating Brownian control problem using formal limiting arguments in the spirit of Harrison [17, 18]. Under a complete resource pooling assumption, the Brownian control problem is reduced to a one-dimension equivalent workload formulation, which is then solved analytically. However, the assumption of a single infinite-server node effectively enforces homogeneous travel times and origin-independent destination routing. This modeling simplification was necessary to permit a one-dimensional state space collapse, but such a reduction is not possible when the model includes multiple infinite-server nodes.
@@
In contrast, our model accommodates multiple infinite-server stations and a two-level routing structure that allows for heterogeneous travel times and captures general origin-destination routing distributions. (Braverman et al. [11] considers a related model with empty-car routing control but studies it under fluid scaling, which avoids the complexities associated with diffusion-scaled fluctuations.) Our setting leads to a multidimensional diffusion limit. By rigorously proving a weak convergence result in this setting, our work provides the theoretical foundation for studying more complex control policies and generalizes the modeling framework introduced in Alwan et al. [1].
@@
As mentioned earlier, the presence of multiple infinite-server stations does not permit a one-dimensional state space collapse, implying that the limiting diffusion control problem remains high-dimensional. However, recent advancements in computing power have renewed interest in studying high-dimensional diffusion control problems using approximate dynamic programming and neural network-based methods; see, e.g., Ata and Kaşıkaralar [5] for an application to a telephone call center. One expects the dynamic control problem associated with the system here to involve drift and singular control of (reflecting) Brownian motion in the heavy traffic limit. Such problems have been sucessfully addressed in dimensions up to 100 or more in recent work by Ata et al. [2, 4], Ata and Xu [6], and Ata and Zhou [7]. While those papers consider different applications, we expect their computational methods to be applicable to ride-hailing applications as well. Although the study of dynamic control of ride-hailing systems in high dimensions is left for future work, our paper lays the mathematical groundwork for future research in stochastic control of such systems, including those where dimensionality precludes analytical tractability.
@@
The rest of the paper is organized as follows. Section 2 introduces the model primitives and makes a sample-path construction of the queue length processes describing the network. Section 3 articulates the heavy traffic assumption and the asymptotic regime, and states the main result of the paper (Theorem 1). Section 4 develops the key tools needed to prove the main result. Section 5 is devoted to a proof of the main result, which involves proving weak convergence of the fluid-scaled processes. Relevant notation and technical preliminaries are given shortly. Additional proofs are provided in Appendices A, B, and C.
@@
For $x,y\in D$ and $T>0$, the following inequalities holds:
$\|\psi(x)\,-\,\psi(y)\|_{T}$ $\,\leq\,\|x\,-\,y\|_{T},$ (3)
$\|\phi(x)\,-\,\phi(y)\|_{T}$  $\,\leq\,2\|x\,-\,y\|_{T};$ (4)
see, e.g., Whitt [26, Lemma 13.5.1]). The space $D^{k}$ is a topological space when endowed with the Skorokhod $J_{1}$ topology. In this topology, a sequence $\left\{x^{n}\right\}_{n=1}^{\infty}$ in $D^{k}$ converges to an element $x\in D^{k}$, written $x^{n}\rightarrow x$, as $n\rightarrow\infty$, if for all continuity points $T>0$ of $x$,
$d_{T}^{k}\big(x^{n}|_{[0,T]},x|_{[0,T]}\big)\,\rightarrow\,0\quad\text{as}\quad n\rightarrow\infty,$
where $d_{T}^{k}:D_{T}^{k}\times D_{T}^{k}\rightarrow[0,\infty)$ is given by
where $e:[0,T]\rightarrow[0,T]$ is the identity map, i.e., $e(t)=t$ for $t\in[0,T]$, and
$\Lambda_{T}\coloneqq\big\{\lambda$ $:[0,T]\rightarrow[0,T]$
$|\,\lambda\text{ is an increasing homeomorphism}\big\}.$
All random variables in this paper are defined on a common probability space $(\Omega,\mathcal{F},P)$. We denote by $\mathcal{M}^{k}$ the Borel $\sigma$-algebra on $D^{k}$ induced by the Skorokhod $J_{1}$ topology. All stochastic processes in this paper are measurable functions from $(\Omega,\mathcal{F},P)$ to $(D^{k},\mathcal{M}^{k})$ for some appropriate dimension $k$. For a sequence of stochastic processes $\{\xi^{n}\}_{n=1}^{\infty}$ and a stochastic process $\xi$, all with sample paths in $D^{k}$ almost surely, we write $\xi^{n}\Rightarrow\xi$ as $n\rightarrow\infty$ to mean that the sequence of probability measures on $(D^{k},\mathcal{M}^{k})$ induced by the processes $\xi^{n}$ converge weakly to the probability measure on $(D^{k},\mathcal{M}^{k})$ induced by the stochastic process $\xi$; see Billingsley [8] and Whitt [26] for further details.
@@ 2.1 Model Primitives
The stochastic evolution of the system is driven by the following primitives. For each $j\in[J]$, let $N_{j}=\{N_{j}(t):t\geq 0\}$ be a unit-rate Poisson process governing service completions at single-server station $j$. For each $k\in[K]$, let $M_{k}=\{M_{k}(t):t\geq 0\}$ be a unit-rate Poisson process governing service completions at infinite-server station $k$. To model the probabilistic routing of jobs in the network, for $j\in[J]$ and $k\in[K]$ we let
$\phi_{j}\,=\,\{\phi_{j}(l):l\geq 1\},\quad\psi_{k}\,=\,\{\psi_{k}(l):l\geq 1\}$
denote sequences of i.i.d. random (routing) vectors. Their probability distributions are given by
@@
To track the cumulative routing of jobs, for each $j\in[J]$ and $k\in[K]$, we define the cumulative routing processes $\Phi_{jk}=\{\Phi_{jk}(m):m\geq 1\}$ and $\Psi_{kj}=\{\Psi_{kj}(m):m\geq 1\}$ as follows:
@@ 2.2 State Dynamics
where $\Phi_{jk}$ and $\Psi_{kj}$ are given by (6), and $T_{j}=\{T_{j}(t):t\geq 0\}$ is a stochastic process that represents the scheduling policy for the server at the $j$th single-server station. To be more specific, $T_{j}(t)$ is the cumulative amount of time the server is working up to time $t$ at single-server station $j$. The corresponding idleness process $I_{j}=\{I_{j}(t):t\geq 0\}$ for $j\in[J]$ of the server at single-server station $j$ is defined as follows:
@@
Equation (15) reflects the obvious physical restriction that the servers can only work on jobs when the queues are not empty. Moreover, (16) implies that the server idleness at each single-server station does not increase as long as the queue is not empty. In other words, we restrict attention to (otherwise arbitrary) work-conserving scheduling policies. Finally, (17)–(18) are natural consequences of the interpretation of $I_{j}$ as server idleness and are standard in the heavy traffic literature. [^1] In the context of ridesharing, this class of scheduling policies can be interpreted as a closest driver policy where a customer is immediately assigned to an available driver in the same region.
@@ 3 Heavy Traffic Assumption and Main Result
Roughly speaking, the condition (20) assumes that every single-server is fully utilized, whereas the condition (21) corresponds to assuming almost all jobs are in the infinite-server stations; see Alwan et al. [1] and Ata et al. [3] for similar assumptions in the context of ride-hailing and volunteer engagement applications, respectively. [^2] This assumption can easily be generalized to limit conditions as follows. Suppose that $\mu_{j}$ depends on $n$, for which we write $\mu_{j}(n)$, and that $\mu_{j}(n)\rightarrow\mu_{j}\in\mathbb{R}$ as $n\rightarrow\infty$ for each $j\in[J]$. Then (20) can be replaced by $\sqrt{n}\,\big[\sum_{k=1}^{K}\sum_{i=1}^{J}\mu_{i}(n)p_{ik}q_{kj}-\mu_{j}(n)\big]\rightarrow c_{j}\in\mathbb{R}$ as $n\rightarrow\infty$ for all $j\in[J]$. Our main result (Theorem 1) still holds, except that we have an additional drift term $c_{j}$. A similar limit condition can be used in place of (21).
To shed light on (20), observe that the expression $\sum_{i=1}^{J}n\mu_{i}p_{ik}$ represents the total rate of jobs leaving the single-server stations to infinite-server station $k$. Therefore, the expression $\sum_{k=1}^{K}q_{kj}\sum_{i=1}^{J}n\mu_{i}p_{ik}$ represents the total rate of jobs entering single-server station $j$ from the infinite-server stations. Because $n\mu_{j}$ is the service rate of the server in single-server station $j$, (20) states that the rate of jobs entering the single-server stations are balanced out by the service rates at the stations.
To shed light on (21), again note that the expression $n\sum_{j=1}^{J}\mu_{j}p_{jk}$ represents the arrival rate to the $k$th infinite-server station, whereas its service rate is $\eta_{k}$. Based on intuition from the classical $M/M/\infty$ queue, we expect the steady-state average queue length at the $k$th infinite-server to be $n\sum_{j=1}^{J}\mu_{j}p_{jk}/\eta_{k}$. It follows that the expected fraction of jobs at the $k$th infinite-server is $m_{k}$. Therefore, (21) states that almost all jobs are at the infinite-server stations. This is consistent with the first condition: It is well known that in large balanced-flow systems in heavy traffic, the queue length processes are of second-order relative to the system size. Consequently, the total number of jobs in the single-server stations are of order $\sqrt{n}$, which implies that the total number of jobs at the infinite-server stations are of order $n$, since the network is closed.
@@
By (7)–(8), (22)–(34) and Assumption 1, it is straightforward verify that for $j\in[J]$, $k\in[K]$, and all $t\geq 0$ the following equalities hold:

### v3

@@ 1 Introduction
The model was motivated in part by ride-hailing systems such as Uber and Lyft. In this context, the single-server stations represent regions of a city where drivers await ride requests, while the infinite-server stations capture travel times between regions. The service rates at the single-server stations represent customer arrival rates in each region, while the service rates at the infinite-server stations represent the travel time completion rates. After picking up a passenger (i.e., after a service completion at a single-server station), a driver enters a travel phase to reach the destination, and upon completion (i.e., after a service completion at an infinite-server station), becomes available again in a potentially new region. After picking up a passenger, a driver enters a travel phase to reach the destination, and upon completion, becomes available again in a potentially new region. The two-level routing structure in our model naturally reflects such origin-destination dynamics, and the presence of multiple infinite-server stations allows for heterogeneity in travel times across different region pairs. While ride-hailing serves as a useful motivation, the model and results apply more broadly to systems in which tasks are queued for assignment or dispatch at specific locations, and then processed or delayed in parallel elsewhere. By establishing a rigorous diffusion approximation for such systems, this paper lays the theoretical groundwork for future research on their dynamic control and optimization.
Ride-hailing systems are often modeled as open, with drivers entering and leaving the platform and hence a time-varying fleet size; for example, Özkan and Ward [22] models a ride-hailing system as an open two-sided matching system. Our model is closed and is most appropriate over time horizons in which the fleet size is approximately constant; over longer horizons, the diffusion approximation can be applied in a piecewise fashion by updating the fleet size between intervals. Demand can also be non-stationary, particularly in the presence of dynamic pricing. Although this paper focuses on time-homogeneous primitives, such non-stationarity can be captured through time-varying service rates at the single-server stations, and we expect the convergence results to extend to such a time-inhomogeneous setting, although the mathematical details are beyond the scope of this paper. In particular, second-order adjustments to the effective demand rates would appear as drift terms in the limit and naturally give rise to drift-rate control problems; see, e.g., Alwan et al. [1]. Extending the convergence results to open networks with external arrivals and departures, while allowing jobs to recirculate through the network, together with infinite-server travel-time nodes, is left for future work.
@@ 1.1 Literature Review
This paper contributes to the literature on heavy traffic diffusion approximations for queueing networks, with a particular focus on closed systems involving infinite-server stations. Classical work by Iglehart [17], Borovkov [8], and Whitt and Glynn [26] establishes diffusion limits for open queueing networks with infinite-server queues. Other papers have studied the structure and properties of these limiting processes. For example, Glynn [13] establishes necessary and sufficient conditions under which the heavy-traffic limit of the $GI/G/\infty$ queue is Markovian. Foundational work on heavy-traffic approximations for stochastic flow networks with bottlenecks includes Chen and Mandelbaum [10, 11], which develop both fluid and diffusion limits. Various surveys, such as Glynn [14] and Pang et al. [23] discuss a variety of techniques for deriving diffusion approximations in queueing systems; see the references therein for a broader overview.
@@
In the context of closed queueing networks with infinite-server stations, several papers establish diffusion limits under heavy traffic conditions. Notably, Kogan et al. [20], Krichagina [21], and Kogan and Lipster [19] analyze models with many single-server stations and a single infinite-server station. These papers have a single probability routing vector and use martingale methods to prove convergence. Our paper supplements this line of research by allowing an arbitrary number of infinite-server stations and a two-level probabilistic routing structure. (This enables more realistic modeling of origin-destination dynamics and heterogeneous travel times, which are important to service systems such as ride-hailing.) Our analysis also differs methodologically: we use a continuous mapping approach and establish existence, uniqueness, and continuity of a nonlinear regulator mapping—results that are not readily implied by existing work. To the best of our knowledge, this is the first paper to rigorously establish a diffusion limit for a closed queueing network with both multiple infinite-server stations and a two-level routing structure.
We note that the network studied here is a closed queueing network with Markovian routing and service times (i.e., a Gordon–Newell network) with a fixed number of jobs $n$, and therefore admits a product-form stationary distribution. Using this stationary distribution, however, requires computing the normalizing constant, i.e., the partition function $G(n)$, which is a sum over all ways of allocating the $n$ jobs across stations; the number of such allocations grows combinatorially with the number of jobs and stations. This quickly becomes computationally difficult to evaluate in large networks. While this challenge has motivated work on asymptotic and algorithmic methods for approximating $G(n)$ in large closed networks, exact computation can still be burdensome when both the number of jobs and the number of stations are large; see, e.g., Birman and Kogan [7] and Kogan [18]. In contrast, the diffusion limit developed here provides a tractable process-level approximation for $\sqrt{n}$-scale fluctuations, capturing transient behavior and serving as a natural basis for diffusion-scale control and optimization in heavy traffic.
@@
A closely related and important antecedent of our model is the one developed in Alwan et al. [1], which considers a ride-hailing system modeled as a closed queueing network with multiple single-server regions and a single infinite-server travel node. The objective is to maximize system profit by making dynamic pricing and dispatch control decisions. That paper formulates an approximating Brownian control problem (BCP) using formal limiting arguments in the spirit of Harrison [15, 16]. Under a complete resource pooling assumption, the BCP is reduced to a one-dimensional equivalent workload formulation, which is then solved analytically. However, the assumption of a single infinite-server node effectively enforces homogeneous travel times and origin-independent destination routing. This modeling simplification was necessary to permit a one-dimensional state space collapse, but such a reduction is not possible when the model includes multiple infinite-server nodes.
@@
In contrast, our model accommodates multiple infinite-server stations and a two-level routing structure that allows for heterogeneous travel times and captures general origin-destination routing distributions. (Notably, Braverman et al. [9] considers a related model with empty-car routing control but studies it under fluid scaling, which avoids the complexities associated with diffusion-scaled fluctuations.) Our setting leads to a multidimensional diffusion limit. By rigorously proving a weak convergence result in this setting, our work provides the theoretical foundation for studying more complex control policies and generalizes the modeling framework introduced in Alwan et al. [1].
@@
As mentioned earlier, the presence of multiple infinite-server stations does not permit a one-dimensional state space collapse, implying that the limiting diffusion control problem remains high-dimensional. However, recent advancements in computing power have renewed interest in studying high-dimensional diffusion control problems using approximate dynamic programming and neural network-based methods. One expects the dynamic control problem associated with the system here to involve drift and singular control of (reflecting) Brownian motion in the heavy-traffic limit. Such problems have been successfully addressed in dimensions up to 100 or more in recent work by Ata et al. [2, 4], and Ata and Xu [5]. While those papers consider different applications, we expect their computational methods to be applicable to ride-hailing applications as well. Although the study of dynamic control of ride-hailing systems in high dimensions is left for future work, our paper lays the mathematical groundwork for future research in stochastic control of such systems, including those where dimensionality precludes analytical tractability.
@@
The rest of the paper is organized as follows. Section 2 introduces the model primitives and makes a sample-path construction of the queue-length processes describing the network. Section 3 articulates the heavy traffic assumption and the asymptotic regime, and states the main result of the paper (Theorem 1). Section 4 develops the key tools needed to prove the main result. Section 5 is devoted to a proof of the main result, which involves proving weak convergence of the fluid-scaled processes. Relevant notation and technical preliminaries are given shortly. Additional proofs are provided in Appendices A, B, C, and D.
@@
The space $D^{k}$ is a topological space when endowed with the Skorokhod $J_{1}$ topology. All random variables in this paper are defined on a common probability space $(\Omega,\mathcal{F},P)$. We denote by $\mathcal{M}^{k}$ the Borel $\sigma$-algebra on $D^{k}$ induced by the Skorokhod $J_{1}$ topology. All stochastic processes in this paper are measurable functions from $(\Omega,\mathcal{F},P)$ to $(D^{k},\mathcal{M}^{k})$ for some appropriate dimension $k$. For a sequence of stochastic processes $\{\xi^{n}\}_{n=1}^{\infty}$ and a stochastic process $\xi$, all with sample paths in $D^{k}$ almost surely, we write $\xi^{n}\Rightarrow\xi$ as $n\rightarrow\infty$ to mean that the sequence of probability measures on $(D^{k},\mathcal{M}^{k})$ induced by the processes $\xi^{n}$ converge weakly to the probability measure on $(D^{k},\mathcal{M}^{k})$ induced by the stochastic process $\xi$; see Billingsley [6] and Whitt [25] for further details.
@@ 2.1 Model Primitives
The stochastic evolution of the system is driven by the following primitives. For $j\in[J]$, let $N_{j}=\{N_{j}(t):t\geq 0\}$ be a unit-rate Poisson process governing service completions at single-server station $j$. For $k\in[K]$, let $M_{k}=\{M_{k}(t):t\geq 0\}$ be a unit-rate Poisson process governing service completions at infinite-server station $k$. To model the probabilistic routing of jobs in the network, let $\phi_{j}=\{\phi_{j}(l):l\geq 1\}$ for $j\in[J]$ and $\psi_{k}=\{\psi_{k}(l):l\geq 1\}$ for $k\in[K]$ denote sequences of i.i.d. random (routing) vectors. Their probability distributions are given by
@@
To track the cumulative routing of jobs, for $j\in[J]$ and $k\in[K]$, we define the cumulative routing processes $\Phi_{jk}=\{\Phi_{jk}(m):m\geq 1\}$ and $\Psi_{kj}=\{\Psi_{kj}(m):m\geq 1\}$ as follows:
@@ 2.2 State Dynamics
where $\Phi_{jk}$ and $\Psi_{kj}$ are given by (4), and $T_{j}=\{T_{j}(t):t\geq 0\}$ is a process that represents the scheduling policy for the server at the $j$th single-server station. To be more specific, $T_{j}(t)$ is the cumulative amount of time the server is working up to time $t$ at single-server station $j$. The corresponding idleness process $I_{j}=\{I_{j}(t):t\geq 0\}$ for $j\in[J]$ of the server at single-server station $j$ is defined as follows:
@@
Equation (13) reflects the obvious physical restriction that the servers can only work on jobs when the queues are not empty. Equation (14) implies that the server idleness at each single-server station does not increase as long as the queue is not empty. In other words, we restrict attention to (otherwise arbitrary) work-conserving scheduling policies. Finally, (15)–(16) are natural consequences of the interpretation of $I_{j}$ as server idleness and are standard in the heavy traffic literature.
@@ 3 Heavy Traffic Assumption and Main Result
Roughly speaking, condition (18) assumes that every single-server is fully utilized, whereas (19) roughly says that almost all jobs are in the infinite-server stations; see Alwan et al. [1] and Ata et al. [3] for similar assumptions in ride-hailing and volunteer engagement applications, respectively. For simplicity, Assumption 1 focuses on the zero-drift case. However, this assumption can be refined to give a nonzero drift term in the heavy-traffic limit. In particular, suppose that the service rates at the single-server stations in the sequence satisfy
$\sqrt{n}\,\Big[\sum_{k=1}^{K}\sum_{i=1}^{J}\mu_{i}p_{ik}q_{kj}\,-\,n^{-1}\mu_{j}^{n}\Big]\rightarrow c_{j}\in\mathbb{R}\quad\text{as}\quad n\rightarrow\infty,$
for all $j\in[J]$. Under this condition, our main result, i.e., Theorem 1, remains valid, except that the limiting diffusion for the single-server queue-length process acquires an additional drift term $c_{j}t$.
To shed light on (18), note that $\sum_{i=1}^{J}n\mu_{i}p_{ik}$ represents the total rate of jobs leaving the single-server stations to infinite-server station $k$. Thus, $\sum_{k=1}^{K}q_{kj}\sum_{i=1}^{J}n\mu_{i}p_{ik}$ represents the total rate of jobs entering single-server station $j$ from the infinite-server stations. Because $n\mu_{j}$ is the service rate of the server in single-server station $j$, (18) states that the rate of jobs entering the single-server stations are balanced out by the service rates at the stations.
To shed light on (19), note that $n\sum_{j=1}^{J}\mu_{j}p_{jk}$ represents the arrival rate to the $k$th infinite-server station, whereas its service rate is $\eta_{k}$. Based on intuition from the classical $M/M/\infty$ queue, we expect the steady-state average queue length at the $k$th infinite-server to be $n\sum_{j=1}^{J}\mu_{j}p_{jk}/\eta_{k}$. It follows that the expected fraction of jobs at the $k$th infinite-server is $m_{k}$. Therefore, (19) states that almost all jobs are at the infinite-server stations. This is consistent with the first condition: It is well known that in large balanced-flow systems in heavy traffic, the queue-length processes are of second-order relative to the system size. Thus, the total number of jobs in the single-server stations are of order $\sqrt{n}$, which implies that the total number of jobs at the infinite-server stations are of order $n$, since the network is closed.
@@
By (5)–(6), (20)–(32), and Assumption 1, it is straightforward verify that for $j\in[J]$, $k\in[K]$, and all $t\geq 0$ the following equalities hold: